P6.1A (LO 1), AN Midlands Inc. had a bad year in 2019. For the first time in its history, it operated at a loss. The companys income statement showed the following results from selling 80,000 units of product: net sales $2,000,000; total costs and expenses $2.235,000; and net loss $235,000. Costs and expenses consisted of the following. Total Variable Fixed Cost of goods sold Selling expenses Administrative expenses $1,568,000 517,000 150,000 $2,235,000 $1.050,000 92,000 58,000 1,200,000 $ 518,000 425,000 92,000 $1,035,000 Management is considering the following independent alternatives for 2020 I. Increase unit selling price 25% with no change in costs and expenses. 2. Change the compensation of salespersons from fixed annual salaries totaling $200,000 to total salaries of $40,000 plus a 5% commission on net sales. 3. Purchase new high-tech factory machinery that will change the proportion between variable and fixed cost of goods sold to 50:50 Instructions a. Compute the break-even point in dollars for 2019 b. Compute the break-even point in dollars under each of the alternative courses of action for 2020 (Round to the nearest dollar.) Which course of action do you recommend?
Answer #1
per unit 80000 units
Sales 25 2000000
Less: Variable costs
Cost of goods sold 13.125 1050000
Selling expenses 1.15 92000
Administrative expenses 0.725 58000
Total variable costs 15 1200000
Contribution 10 800000
CM Ratio(Unit Contribution/Unit Price) 40%
a.
Break-even point in dollars for 2019 = Total Fixed cost/CM Ratio
=1035000/40%= 2587500
b.
1
per unit
Sales 31.25
Less: Variable costs
Cost of goods sold 13.125
Selling expenses 1.15
Administrative expenses 0.725
Total variable costs 15
Contribution 16.25
CM Ratio(Unit Contribution/Unit Price) 52%
Break-even point in dollars for 2020 = Total Fixed cost/CM Ratio
=1035000/52%= 1990385
2
Sales 25
Less: Variable costs
Cost of goods sold 13.125
Selling expenses 1.15
Administrative expenses 0.725
Commission 1.25
Total variable costs 16.25
Contribution 8.75
CM Ratio(Unit Contribution/Unit Price) 35%
Fixed costs = 1035000-(200000-40000)= 875000
Break-even point in dollars for 2020 = Total Fixed cost/CM Ratio
=875000/35%= 2500000
3
Sales 25
Less: Variable costs
Cost of goods sold 9.8 =(1568000/2)/80000
Selling expenses 1.15
Administrative expenses 0.725
Total variable costs 11.675
Contribution 13.325
CM Ratio(Unit Contribution/Unit Price) 53.30%
Fixed costs=1035000+(1568000/2-518000)= 1301000
Break-even point in dollars for 2020 = Total Fixed cost/CM Ratio
=1301000/53.3%= 2440901
Alternative 1 is the best as it has least break even Dollar sales.